OpenStudy (anonymous):

May you check my answers? I got D. and I think the last one looks like this: 4+6= 2(2+3) @ganeshie8

OpenStudy (anonymous):

A and D

OpenStudy (anonymous):

@ganeshie8 Please help me..

OpenStudy (anonymous):

Any doubt??

OpenStudy (anonymous):

@waterineyes Yes!!!! I need help with both of them. I got option D for the first one and I think the last one looks like this: 4+6= 2(2+3)

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Yes for last one you are very well right.. 4 + 6 = 2(2 + 3), so here what property you are using??

OpenStudy (anonymous):

There?? Tell me here which property is being used??

OpenStudy (anonymous):

Is the first question option D or A? I think the last one is distributive.

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Yes the last one is Distributive.. \(a(b + c) = ab + ac\) or \(ab + ac = a(b+c)\) This property is Distributive Property.. So, for last one D is right.

OpenStudy (anonymous):

But for first one, how are you getting D?? Can you explain, where is the doubt coming??

OpenStudy (anonymous):

Well you are distributing the 3, correct? I'm confused because some people have told me it's associative! :(

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Why are you replying so late??

OpenStudy (anonymous):

There is no Distributive Property in Multiplication case..

OpenStudy (anonymous):

\[a(bc ) = (ab)c\] It is Associative Property of Multiplication..

OpenStudy (anonymous):

Here, you are not distributing a to b and c, you are just changing the brackets.. Earlier, brackets were with bc and now brackets are with ab..

OpenStudy (anonymous):

Similarly, \(3(3 \cdot y) = (3 \cdot 3 ) y\)

OpenStudy (anonymous):

You must know it : \(a + b = b + a\) \(\implies \text{Commutative Property of Addition}\) \(a + (b + c) = (a + b) + c\) \(\implies \text{Associative Property of Addition}\) \(a(b + c) = ab + ac\) \(\implies \text{Distributive Property of Addition}\) \(ab = ba\) \(\implies \text{Commutative Property of Multiplication}\) \(a(bc) = (ab)c\) \(\implies \text{Associative Property of Multiplication}\)

OpenStudy (anonymous):

Thank you so much, sorry for the delay- OS doesn't send me the notifications :/

OpenStudy (anonymous):

Okay, then you must stick to your post... :P

OpenStudy (anonymous):

:)